OpenStudy (anonymous):

Describe the roots of the equation shown below. 3x^2+4x+2=0

OpenStudy (anonymous):

There are two real, rational roots. There are two real, irrational roots. There is one real, double root. There are two complex roots.

OpenStudy (***[isuru]***):

find the discriminant of the above equation... let's compare it with our model.. ax^2 + bx + c = 0 3x^2 + 4x + 2=0 here in ur problem... a = 3 b = 4 c = 2 so.... the discriminant is b^2 - 4ac which is 4^2 - 4(2*3) = 16 - 24 = -8 so the discriminant is negative.. which means there r 2 complex roots.....

OpenStudy (anonymous):

thanks!

OpenStudy (***[isuru]***):

u r welcome!!

OpenStudy (wolf1728):

Just thought I'd add the reason why there are 2 complex roots. You have to take the square root of the discriminant (b² - 4ac). When b² -4ac is negative, you have to determine the square root of a negative number which always results in a complex number.